OpenStudy (anonymous):

Solve & Check 5x + 7 = 2x - 14

OpenStudy (anonymous):

1) subtract \(2x\) from both sides 2) subtract 7 from both sides 3) divide both sides by 3

OpenStudy (anonymous):

@satellite73 I got 3x = -7. so i divided & got x= 2 1/7. Now i"m confused on the check part

OpenStudy (anonymous):

Now that you know the value of x (x=2 1/7), plug 2 1/7 for x back into the original equation 5x + 7 = 2x - 14. See if both side are equal to one another.

OpenStudy (anonymous):

when you subtract \(7\) from both sides you should get \(3x=-21\) because \(-14-7=-21\)

OpenStudy (anonymous):

lets work through the steps \[5x + 7 = 2x - 14 \] \[3x+7=-14\] \[3x=-21\] \[x=-\frac{21}{3}=-7\]

OpenStudy (anonymous):

it is a common mistake to drop the minus sign when the term in front is gone, but remember the right hand side had \(2x-14\) so when you subtract \(2x\) you goet \(-14\) not \(14\) the minus sign comes with the number

OpenStudy (anonymous):

@satellite73 Thank You SOOO much!!

OpenStudy (anonymous):

yw my pleasure